OpenStudy (anonymous):

x^4 - 8x^2 = -1 is my first step to square both sides? if so then whats' the next step?

geerky42 (geerky42):

Let x² = u; u² - 8u = -1 Now add both sides by 16. u² - 8u + 16 = 15 (u - 4)² = 15 u - 4 = ±√15 u = 4 ± √15 x² = 4 ± √15 \(x = \pm\sqrt{4\pm\sqrt{15}}\) \(\boxed{x = \sqrt{4+\sqrt{15}}~\text{or}~\sqrt{4-\sqrt{15}}~\text{or}~-\sqrt{4+\sqrt{15}}~\text{or}~-\sqrt{4-\sqrt{15}}}\)

OpenStudy (anonymous):

oh okay, i would not have thought of that one. thanks